# A Systematic Literature Review of Spatio-Temporal Graph Neural Network Models for Time Series Forecasting and Classification

### v2

@@ 6 Discussion
Energy 1.6 5.8 3%
Environment 1.3 6.5 26%
Finance 1.8 5.6 12%
Generic 4.2 10.0 41%
Health 1.7 7.7 53%
Mobility 2.6 8.4 27%
Other 2.1 5.6 14%
Predictive monitoring 2.1 8.1 14%
@@ 7 Limits, challenges, and future research directions
Comparability. As highlighted in Ref. [cit_rev2], the evaluation of GNN models has improved thanks to the introduction of the Open Graph Benchmark (OGB) [NEURIPS2020_fb60d411], which provides a standardized evaluation framework and a variety of benchmark graph datasets. However, there is currently no standardized benchmark for spatio-temporal GNNs. As a result, each model is evaluated on its own selection of datasets, and this fragmentation makes it difficult to compare the results of different studies. To alleviate this problem, this review presented all the information gathered from the selected papers, including datasets, benchmarks, codes, and tables of results. All that is intended to serve as a detailed overview and a foundation for further exploration, in the hope that researchers will begin to examine the data collected in the presented tables and, over time, identify relevant datasets and benchmark models. Moreover, we encourage the authors to discuss the relationship between complexity and performance of the proposed models to facilitate a more comprehensive comparison.
@@
Explainability. A crucial aspect for GNN models is explainability, intended as the ability to interpret and understand the decision-making processes implemented in them. Despite their attempt to explicitly model spatial relationships between series, GNNs are often considered as “black box” models. Remarkably, none of the selected papers adequately tackle the concept of explainability. Some models are claimed to be explainable because of the integration of non-linear physical laws (e.g., Ref. [art_401]), while other papers claim explainability on the basis of the possibility to visualize a graph (e.g., Ref. [art_360]) or because they present a feature aggregation form that can be intuitively understood. In addition, other approaches focus on determining feature importance through tools such as Captum [Captum], which enables the interpretation and understanding of PyTorch model predictions by exploring the features that contribute to a specific outcome (e.g., Ref. [art_361]). This highlights a research gap in this area. However, recent literature is beginning to investigate this important topic. For instance, few papers studied perturbation-based explanation methods, which however have proven to be somewhat ineffective, as reported in Ref. [interpret_3]. Other approaches, such as GNNExplainer [interpret_2], aim to identify a small subset of node features that have a crucial role in GNN’s predictions, as in Ref. [art_349, art_357]. Recently, some researchers have started to explore graph counterfactuals as a means of generating explanations, as reviewed in Ref. [interpret_1].
@@
Quasi-equilibrium learning. As pointed out in Subsec. 4.2, GNNs were originally introduced in Ref. [Scarselli2009] by promoting an extension of the backpropagation algorithm which exploits a fixed-point principle. The subsequent introduction of the notion of convolutional graph has fundamentally changed such computational scheme. It is worth mentioning that in contemporary GNNs there is no need to wait for the system to reach equilibrium. For example, the deep neural network scheme no longer requires weight-sharing. These improvements have led to successful applications in many real-world problems, where modern features contribute significantly to both accuracy and computational efficiency. However, when analyzing the processing of temporal information, GNNs are often implemented within a single graph domain where a temporal function is defined. This is a case where the original computational approach proposed in Ref. [Scarselli2009] seems to be more adequate, especially when considering that the rate of weight updates is generally much slower than the rate of external temporal signals. This suggests the necessity of revisiting the process of convergence to the fixed point, and of exploring computational schemes which take place at quasi-equilibrium points. We expect that this approach can significantly alleviate the computational burden that is typical of the convergence to fixed-points required for spatio-temporal graph collections.
@@ 8 Conclusions
Another key point which results from the presented overview is that the current literature on spatio-temporal GNNs is very fragmented, as it was pictorially reported in Fig. 3. This can be attributed to the fact that the involved researchers come from different communities which focus on distinct application domains. As a result, there seems to be a lack of standardized datasets or benchmarks in the studies. The objective of this review was also to collect information on datasets, proposed models, links to source codes, benchmarks, and results, in order to provide common background information for future studies.
@@
Link to the GitHub repository with the links to the 281 reviewed papers, organized by groups: https://github.com/FlaGer99/SLR-Spatio-Temporal-GNN.git.
@@
Flavio Corradini: Project administration, Funding acquisition. Flavio Gerosa: Investigation, Writing - Review & Editing, Visualization. Marco Gori: Writing - Review & Editing, Supervision. Carlo Lucheroni: Writing - Review & Editing, Supervision. Marco Piangerelli: Conceptualization, Methodology, Writing - Review & Editing, Supervision. Martina Zannotti: Conceptualization, Methodology, Validation, Investigation, Writing - Original Draft, Writing - Review & Editing, Visualization.
@@ Acknowledgments
Appendix A List of selected journal papers
@@
Tab. LABEL:t:selected_papers lists all the journal papers included in this SLR, together with the year of publication, group they belong to, case study, and nature of the task (classification or forecasting).
@@
Table 41: List of the 281 selected journal papers, with year, group they belong to, case study, and nature of the task.
Ref. Year Group Case study Task
[art_15] 2023 Energy Heat load Forecasting
[art_16] 2023 Energy Electricity net load Forecasting
[art_17] 2023 Energy Residential load Forecasting
[art_18] 2023 Energy Electricity load Forecasting
[art_19] 2023 Energy Photovoltaic power Forecasting
[art_20] 2022 Energy Wind power Forecasting
[art_21] 2022 Energy Photovoltaic power Forecasting
[art_22] 2022 Energy Photovoltaic power Forecasting
[art_23] 2022 Energy Power system transient dynamics Forecasting
[art_24] 2020 Energy Wind power Forecasting
[art_25] 2022 Energy Urban energy consumption Forecasting
[art_26] 2023 Energy Solar power Forecasting
[art_27] 2022 Energy Wind power Forecasting
[art_332] 2024 Energy Wind power Forecasting
[art_333] 2024 Energy Wind power Forecasting
[art_334] 2024 Energy Air conditioning control Forecasting
[art_335] 2024 Energy Air conditioning control Forecasting
[art_337] 2024 Energy Electrical demand and prices Forecasting
[art_338] 2024 Energy Electric load Forecasting
[art_339] 2024 Energy Electric load Forecasting
[art_340] 2024 Energy Chiller energy consumption Forecasting
[art_341] 2024 Energy Electricity demand Forecasting
[art_342] 2024 Energy Electric power anomaly detection Classification
[art_343] 2024 Energy Electricity demand Forecasting
[art_344] 2024 Energy Electricity consumption Forecasting
[art_345] 2024 Energy District heating networks Forecasting
[art_347] 2024 Energy Power quality disturbances Classification
[art_348] 2024 Energy Residential load Forecasting
[art_349] 2024 Energy Photovoltaic energy Forecasting
[art_350] 2024 Energy Electric vehicle charging stations Forecasting
[art_351] 2024 Energy Natural gas consumption Forecasting
[art_353] 2024 Energy Networked microgrids Forecasting
[art_390] 2024 Energy Multivariate time series Forecasting
[art_1] 2023 Environment Air quality index Forecasting
[art_2] 2023 Environment PM 2.5 concentration Forecasting
[art_3] 2023 Environment PM 2.5 concentration Forecasting
[art_4] 2023 Environment PM 2.5 concentration Forecasting
[art_5] 2023 Environment Air quality index Forecasting
[art_6] 2023 Environment PM 2.5 concentration Forecasting
[art_7] 2023 Environment PM 2.5 concentration Forecasting
[art_8] 2022 Environment PM 2.5 concentration Forecasting
[art_28] 2023 Environment Sea temperature Forecasting
[art_29] 2023 Environment Wind speed Forecasting
[art_30] 2023 Environment Water quality Forecasting
[art_31] 2023 Environment Wind speed Forecasting
[art_32] 2023 Environment Significant wave height Forecasting
[art_33] 2023 Environment Sea temperature Forecasting
[art_34] 2023 Environment Wind speed Forecasting
[art_35] 2023 Environment Sea temperature Forecasting
[art_36] 2023 Environment Sea temperature Forecasting
[art_38] 2022 Environment Frost Both
[art_39] 2022 Environment Sea temperature Forecasting
[art_40] 2021 Environment Wind speed Forecasting
[art_41] 2021 Environment Sea temperature Forecasting
[art_42] 2022 Environment Water quality Forecasting
[art_43] 2023 Environment Groundwater level Forecasting
[art_44] 2023 Environment Wind speed Forecasting
[art_45] 2023 Environment Rainfall Forecasting
[art_46] 2023 Environment Wind speed Forecasting
[art_47] 2023 Environment Sea temperature Forecasting
[art_354] 2024 Environment PM2.5 prediction Forecasting
[art_355] 2024 Environment Sea clutter Forecasting
[art_356] 2024 Environment Earthquakes Forecasting
[art_357] 2024 Environment Plankton dynamics Forecasting
[art_359] 2024 Environment Groundwater data Forecasting
[art_360] 2024 Environment Carbon emission Forecasting
[art_361] 2024 Environment PM2.5 prediction Forecasting
[art_363] 2024 Environment Water demand Forecasting
[art_364] 2024 Environment Water quality and flow rate Forecasting
[art_365] 2024 Environment Weather Forecasting
[art_366] 2024 Environment Wind speed scenario Forecasting
[art_367] 2024 Environment Carbon emission Forecasting
[art_368] 2024 Environment Air quality Forecasting
[art_370] 2024 Environment Temperature Forecasting
[art_371] 2024 Environment Air quality Forecasting
[art_373] 2024 Environment Greenhouse gas concentrations Forecasting
[art_59] 2022 Finance Stock prediction Classification
[art_60] 2023 Finance Stock prediction Classification
[art_62] 2023 Finance Stock prediction Classification
[art_63] 2023 Finance Stock prediction Classification
[art_64] 2022 Finance Investment prediction Forecasting
[art_65] 2023 Finance Stock prediction Classification
[art_66] 2022 Finance Financial prediction Classification
[art_67] 2023 Finance Trading Classification
[art_68] 2023 Finance Sale prediction Forecasting
[art_374] 2024 Finance Oil and metal market Classification
[art_375] 2024 Finance Supply chain demand Forecasting
[art_377] 2024 Finance Stock price Classification
[art_378] 2024 Finance Stock market volatility Forecasting
[art_379] 2024 Finance Agricultural price Forecasting
[art_381] 2024 Finance Cryptocurrencies’ price Forecasting
[art_382] 2024 Finance Stock trend Classification
[art_69] 2023 Generic Batch workloads Forecasting
[art_70] 2023 Generic Node classification for websites Classification
[art_71] 2023 Generic Clinical risk classification Classification
[art_72] 2023 Generic Sensor network Forecasting
[art_73] 2023 Generic Multivariate time series Forecasting
[art_74] 2023 Generic Multivariate time series Forecasting
[art_75] 2023 Generic Multivariate time series Forecasting
[art_76] 2023 Generic Long term series forecasting Forecasting
[art_77] 2023 Generic Multivariate time series Forecasting
[art_78] 2023 Generic Multivariate time series Forecasting
[art_79] 2023 Generic Multivariate time series Forecasting
[art_80] 2023 Generic Multivariate time series Forecasting
[art_81] 2023 Generic Multivariate time series Forecasting
[art_82] 2023 Generic Multivariate long sequences Forecasting
[art_83] 2023 Generic Multivariate time series Forecasting
[art_84] 2022 Generic Multivariate time series Both
[art_85] 2023 Generic Multivariate time series Forecasting
[art_86] 2022 Generic Multivariate time series Forecasting
[art_87] 2022 Generic Multivariate time series Forecasting
[art_88] 2023 Generic Multivariate time series Forecasting
[art_89] 2023 Generic Multivariate time series Forecasting
[art_90] 2022 Generic Multivariate time series Forecasting
[art_160] 2023 Generic Impact of graph construction Forecasting
[art_202] 2021 Generic Generic Both
[art_383] 2024 Generic Multivariate time series Forecasting
[art_384] 2024 Generic Multivariate time series Forecasting
[art_385] 2024 Generic Multivariate time series Classification
[art_386] 2024 Generic Multivariate time series Classification
[art_387] 2024 Generic Multivariate time series Both
[art_388] 2024 Generic Multivariate time series Forecasting
[art_389] 2024 Generic Multivariate time series Classification
[art_393] 2024 Generic Multivariate time series Forecasting
[art_395] 2024 Generic Multivariate time series Classification
[art_396] 2024 Generic Multivariate time series Forecasting
[art_397] 2024 Generic Multivariate time series Forecasting
[art_398] 2024 Generic Multivariate time series Forecasting
[art_400] 2024 Generic Multivariate time series Forecasting
[art_401] 2024 Generic Multivariate time series Forecasting
[art_403] 2024 Generic Multivariate time series Both
[art_404] 2024 Generic Multivariate time series Forecasting
[art_405] 2024 Generic Multivariate time series Forecasting
[art_407] 2024 Generic Multivariate time series Classification
[art_417] 2024 Generic Multivariate time series Forecasting
[art_464] 2024 Generic Multivariate time series Forecasting
[art_14] 2023 Health Epilepsy diagnosis Classification
[art_48] 2023 Health Epidemic prediction Forecasting
[art_49] 2023 Health Epidemic prediction Forecasting
[art_50] 2022 Health Epidemic prediction Forecasting
[art_51] 2021 Health Epidemic prediction Forecasting
[art_52] 2022 Health Epidemic prediction Forecasting
[art_92] 2023 Health Disease diagnosis Classification
[art_189] 2023 Health Clinical risk classification Classification
[art_410] 2024 Health Infectious disease Forecasting
[art_411] 2024 Health EEG emotion recognition Classification
[art_412] 2024 Health EEG emotion recognition Classification
[art_413] 2024 Health Epidemic Forecasting
[art_414] 2024 Health Epidemic Forecasting
[art_415] 2024 Health Epidemic Forecasting
[art_416] 2024 Health Epidemic Forecasting
[art_93] 2022 Mobility Urban traffic Forecasting
[art_94] 2021 Mobility Crowd flow Forecasting
[art_96] 2022 Mobility Crowd flow Forecasting
[art_97] 2023 Mobility Urban traffic Forecasting
[art_98] 2022 Mobility Urban traffic Forecasting
[art_99] 2023 Mobility Urban traffic Forecasting
[art_100] 2023 Mobility Urban traffic Forecasting
[art_102] 2021 Mobility Passenger demand Forecasting
[art_103] 2023 Mobility Urban traffic Forecasting
[art_104] 2022 Mobility Internet of vehicles traffic Forecasting
[art_105] 2020 Mobility Urban traffic Forecasting
[art_106] 2023 Mobility Urban traffic Forecasting
[art_108] 2023 Mobility Urban traffic Forecasting
[art_109] 2023 Mobility Urban traffic Forecasting
[art_110] 2023 Mobility Urban traffic Forecasting
[art_113] 2023 Mobility Urban traffic Forecasting
[art_114] 2023 Mobility Bike station demand Forecasting
[art_115] 2023 Mobility Urban traffic Forecasting
[art_116] 2023 Mobility Urban traffic Forecasting
[art_117] 2023 Mobility Urban traffic Forecasting
[art_119] 2022 Mobility Taxi demand Forecasting
[art_120] 2022 Mobility Urban traffic Forecasting
[art_121] 2022 Mobility Metro passengers Forecasting
[art_122] 2022 Mobility Urban traffic Forecasting
[art_123] 2021 Mobility Urban traffic Forecasting
[art_124] 2022 Mobility Urban traffic Forecasting
[art_125] 2023 Mobility Urban traffic Forecasting
[art_126] 2020 Mobility Urban traffic Forecasting

### v4

@@ 6.2 Observed trends
Energy 1.6 5.7 3%
Environment 1.3 6.6 23%
Finance 1.5 6.3 7%
Generic 4.5 10.6 56%
Health 1.7 8.3 48%
Mobility 2.8 9.4 35%
Other 1.9 6.3 15%
Predictive monitoring 2.2 7.9 22%
6.3 Meta-analysis across groups
Here, we conduct a meta-analysis of the reviewed spatio-temporal GNNs models across multiple application domains. By definition, all models incorporate both spatial and temporal components, either arranged sequentially (space-then-time) or jointly (space-and-time). Beyond this common feature, designs vary widely, and our meta-analysis focuses on the key aspects that characterize these models, namely graph construction and spatial aggregation, aiming to reveal patterns where possible.
However, the highly fragmented research landscape makes identifying these trends challenging. We observed that many models are tailored to specific case studies, and differences in domain focus, limited reproducibility, and inconsistent experimental setups further complicate cross-domain comparisons and the extraction of generalizable insights. For this reason, rather than attempting a meta-analysis across the entire fragmented landscape of reviewed papers, we focused on the best-performing models within each domain to identify shared characteristics and emerging trends. In particular, we selected those models that, in the results tables, ranked among the top performers for the associated datasets and that were briefly discussed below each table. These models were gathered across all groups and are available in the GitHub repository associated with this work.
In this selected subset, models from the “Generic” and “Mobility” thematic groups are more numerous. This is not because they are inherently more important, but because these models offer greater comparability, having been evaluated across multiple common datasets in their original studies. For this reason, they are particularly suitable for this broader meta-analysis, which does not focus on highly specific use cases.
As for the determination of the graph structure, most models in the selected subset have a self-learning module. Pre-defined adjacency matrices based on physical distances, correlations, cosine similarities, or similar techniques remain more common in studies focused on specific tasks. From a historical perspective, early models relied more heavily on pre-defined adjacency matrices, whereas more recent works increasingly favor learned or hybrid adjacency matrices, often combining learned and pre-defined structures to enhance adaptivity and robustness. This shift toward learned or hybrid adjacency matrices, often implemented via self-learning modules, supports broader applicability across diverse domains. Among self-learning modules in spatio-temporal GNNs, two predominant approaches have emerged with comparable prevalence in the selected subset of papers: the method of the GWN model [425] and that of the MTGNN model [424], both developed by Wu et al, and discussed in Subsec. 4.4.
As for the spatial aggregation module, the subset reveals that graph convolution remains the dominant spatial aggregation mechanism in high-performing spatio-temporal GNNs, while attention-based spatial aggregation is less common. This suggests that the practical advantages of convolutional aggregation, in terms of efficiency and training stability, continue to make it preferable to attention mechanisms. The aggregation mechanism of recurrent GNNs, which defined the original GNN formulation, is instead almost entirely unused.
@@ 7 Limits, challenges, and future research directions
Comparability. As highlighted in Ref. [273], the evaluation of GNN models has improved thanks to the introduction of the Open Graph Benchmark (OGB) [146], which provides a standardized evaluation framework and a variety of benchmark graph datasets. However, there is currently no standardized benchmark for spatio-temporal GNNs. As a result, each model is evaluated on its own selection of datasets, and this fragmentation makes it difficult to compare the results of different studies. To alleviate this problem, this review presented all the information gathered from the selected papers, including datasets, benchmarks, codes, and tables of results. All that is intended to serve as a detailed overview and a foundation for further exploration, in the hope that researchers will begin to examine the data collected in the presented tables and, over time, identify relevant datasets and benchmark models. Moreover, we encourage the authors to discuss the relationship between complexity and performance of the proposed models to facilitate a more comprehensive comparison.
@@
Explainability. A crucial aspect for GNN models is explainability, intended as the ability to interpret and understand the decision-making processes implemented in them. Despite their attempt to explicitly model spatial relationships between series, GNNs are often considered as “black box” models. Remarkably, none of the selected papers adequately tackle the concept of explainability. Some models are claimed to be explainable because of the integration of non-linear physical laws (e.g., Ref. [360]), while other papers claim explainability on the basis of the possibility to visualize a graph (e.g., Ref. [106]) or because they present a feature aggregation form that can be intuitively understood. In addition, other approaches focus on determining feature importance through tools such as Captum [192], which enables the interpretation and understanding of PyTorch model predictions by exploring the features that contribute to a specific outcome (e.g., Ref. [325]). This highlights a research gap in this area. However, recent literature is beginning to investigate this important topic, as also discussed in [369]. For instance, few papers studied perturbation-based explanation methods. Other approaches, such as GNNExplainer [455], aim to identify a small subset of node features that have a crucial role in GNN’s predictions, as in Ref. [377, 184]. Recently, some researchers have started to explore graph counterfactuals as a means of generating explanations, as reviewed in Ref. [314]. For a recent survey on GNNs for explainable AI see Ref. [292].
@@
Quasi-equilibrium learning. As pointed out in Subsec. 4.2, GNNs were originally introduced in Ref. [332] by promoting an extension of the backpropagation algorithm which exploits a fixed-point principle. The subsequent introduction of the notion of convolutional graph has fundamentally changed such computational scheme. It is worth mentioning that in contemporary GNNs there is no need to wait for the system to reach equilibrium. For example, the deep neural network scheme no longer requires weight-sharing. These improvements have led to successful applications in many real-world problems, where modern features contribute significantly to both accuracy and computational efficiency. However, when analyzing the processing of temporal information, GNNs are often implemented within a single graph domain where a temporal function is defined. This is a case where the original computational approach proposed in Ref. [332] seems to be more adequate, especially when considering that the rate of weight updates is generally much slower than the rate of external temporal signals. This suggests the necessity of revisiting the process of convergence to the fixed point, and of exploring computational schemes which take place at quasi-equilibrium points. We expect that this approach can significantly alleviate the computational burden that is typical of the convergence to fixed-points required for spatio-temporal graph collections.
Broader Directions. The field of time series modeling by means of GNNs is expected to expand in a set of further directions. While most of the reviewed studies focus on point forecasting with deterministic models, relatively few address probabilistic forecasting (e.g., Refs. [409, 143]), which is crucial for assessing uncertainty in predictions. Another potential direction could be the integration of spatio-temporal GNNs with transformer-based architectures or large language models. Although studies are still limited, such models could help capture long term dependencies and complex patterns in time series data, yet not forgetting all possible limitations of LLMs in time series applications [367]. Exploring these directions may help overcome current limitations in modeling uncertainty and complex temporal relationships, providing more informative and reliable results.
@@ 8 Conclusions
Another key point which results from the presented overview is that the current literature on spatio-temporal GNNs is very fragmented, as it was pictorially reported in Fig. 5. This can be attributed to the fact that the involved researchers come from different communities which focus on distinct application domains. As a result, there seems to be a lack of standardized datasets or benchmarks in the studies. The objective of this review was also to collect information on datasets, proposed models, links to source codes, benchmarks, and results, in order to provide common background information for future studies. To support this goal, the GitHub repository associated with this review offers interactive tools for sorting, filtering, and visualizing the results achieved by the different models, making it easier for researchers to explore and compare existing work.
@@
Link to the GitHub repository containing information on the 366 reviewed papers organized by groups, along with interactive tables and charts: https://github.com/FlaGer99/SLR-Spatio-Temporal-GNN.git.
@@
Flavio Corradini: Project administration, Funding acquisition. Flavio Gerosa: Investigation, Writing - Review & Editing, Visualization. Marco Gori: Writing - Review & Editing, Supervision. Carlo Lucheroni: Writing - Review & Editing, Supervision. Marco Piangerelli: Conceptualization, Methodology, Writing - Original Draft, Writing - Review & Editing, Supervision. Martina Zannotti: Conceptualization, Methodology, Validation, Formal analysis, Investigation, Data Curation, Writing - Original Draft, Writing - Review & Editing, Visualization.
Author names appear in alphabetical order and do not reflect relative contributions.
@@ Acknowledgments
Appendix A List of selected papers
@@
Tab. LABEL:t:selected_papers lists all the papers included in this SLR, together with the year of publication, group they belong to, case study, and nature of the task (classification or forecasting).
@@
Table 41: List of the selected papers, with year, group they belong to, case study, and nature of the task.
Reference Year Group Case study Task
[460] 2020 Energy Wind power Forecasting
[256] 2021 Energy Electric load Forecasting
[132] 2022 Energy Wind power Forecasting
[350] 2022 Energy Photovoltaic power Forecasting
[351] 2022 Energy Photovoltaic power Forecasting
[509] 2022 Energy Power system transient dynamics Forecasting
[147] 2022 Energy Urban energy consumption Forecasting
[216] 2022 Energy Wind power Forecasting
[405] 2023 Energy Heat load Forecasting
[486] 2023 Energy Electricity load Forecasting
[398] 2023 Energy Residential load Forecasting
[93] 2023 Energy Electricity load Forecasting
[317] 2023 Energy Solar power Forecasting
[439] 2024 Energy Wind power Forecasting
[452] 2024 Energy Wind power Forecasting
[531] 2024 Energy Air conditioning control Forecasting
[148] 2024 Energy Air conditioning control Forecasting
[323] 2024 Energy Electrical demand and prices Forecasting
[104] 2024 Energy Electric load Forecasting
[477] 2024 Energy Electric load Forecasting
[265] 2024 Energy Chiller energy consumption Forecasting
[288] 2024 Energy Electricity demand Forecasting
[161] 2024 Energy Electric power anomalies Classification
[324] 2024 Energy Electricity demand Forecasting
[156] 2024 Energy Electricity consumption Forecasting
[18] 2024 Energy District heating networks Forecasting
[24] 2024 Energy Power quality disturbances Classification
[87] 2024 Energy Residential load Forecasting
[377] 2024 Energy Photovoltaic energy Forecasting
[500] 2024 Energy Electric vehicle charging stations Forecasting
[315] 2024 Energy Natural gas consumption Forecasting
[463] 2024 Energy Networked microgrids Forecasting
[319] 2024 Energy Multivariate time series Forecasting
[102] 2021 Environment Wind speed Forecasting
[361] 2021 Environment Sea temperature Forecasting
[42] 2022 Environment Water temperature, stream flow Forecasting
[310] 2022 Environment PM 2.5 concentration Forecasting
[257] 2022 Environment Frost Both
[392] 2022 Environment Sea temperature Forecasting
[5] 2023 Environment Soil Moisture Forecasting
[41] 2023 Environment Air quality index Forecasting
[267] 2023 Environment PM 2.5 concentration Forecasting
[357] 2023 Environment PM 2.5 concentration Forecasting
[223] 2023 Environment PM 2.5 concentration Forecasting
[72] 2023 Environment Air quality index Forecasting
[293] 2023 Environment PM 2.5 concentration Forecasting
[183] 2023 Environment PM 2.5 concentration Forecasting
[437] 2023 Environment Sea temperature Forecasting
[11] 2023 Environment Wind speed Forecasting
[210] 2023 Environment Water quality Forecasting
[326] 2023 Environment Wind speed Forecasting
[502] 2023 Environment Significant wave height Forecasting
[185] 2023 Environment Sea temperature Forecasting
[12] 2023 Environment Wind speed Forecasting
[304] 2023 Environment Sea temperature Forecasting
[98] 2023 Environment Sea temperature Forecasting
[10] 2023 Environment Groundwater level Forecasting
[97] 2023 Environment Wind speed Forecasting
[311] 2023 Environment Rainfall Forecasting
[73] 2023 Environment Wind speed Forecasting
[345] 2023 Environment Sea temperature Forecasting
[133] 2024 Environment Weather Forecasting
[248] 2024 Environment PM2.5 prediction Forecasting
[229] 2024 Environment Sea clutter Forecasting
[184] 2024 Environment Plankton dynamics Forecasting
[366] 2024 Environment Groundwater data Forecasting
[106] 2024 Environment Carbon emission Forecasting
[325] 2024 Environment PM2.5 prediction Forecasting
[416] 2024 Environment Water demand Forecasting
[417] 2024 Environment Water quality and flow rate Forecasting
[336] 2024 Environment Weather Forecasting
[269] 2024 Environment Wind speed scenario Forecasting
[45] 2024 Environment Air quality Forecasting
[403] 2024 Environment Temperature Forecasting
[426] 2024 Environment Air quality Forecasting
[17] 2024 Environment Greenhouse gas concentrations Forecasting
[343] 2022 Finance Energy markets Forecasting
[429] 2022 Finance Stock prices Classification
[151] 2022 Finance Stock prediction Classification
[225] 2022 Finance Investment prediction Forecasting
[53] 2022 Finance Financial prediction Classification
[307] 2023 Finance Stock prediction Classification
[371] 2023 Finance Stock prediction Classification
[391] 2023 Finance Stock prediction Classification
[362] 2023 Finance Stock prediction Classification
[158] 2023 Finance Trading Classification
[444] 2023 Finance Sale prediction Forecasting
[297] 2024 Finance Supply chain demand Forecasting
[205] 2024 Finance Stock market volatility Forecasting
[37] 2024 Finance Stock trend Classification
[68] 2020 Health Epidemics Forecasting
[334] 2021 Health Covid 19 severity Forecasting
[514] 2021 Health Epidemics Forecasting
[290] 2021 Health Epidemic prediction Forecasting
[38] 2022 Health Epilepsy diagnosis Classification
[160] 2022 Health Epidemics Forecasting
[282] 2022 Health Epidemics Forecasting
[235] 2022 Health Epidemic prediction Forecasting
[298] 2022 Health Epidemic prediction Forecasting
[23] 2023 Health Epilepsy diagnosis Classification
[384] 2023 Health Sleep stage, emotion recognition Classification
[383] 2023 Health Epilepsy diagnosis Classification
[483] 2023 Health Epidemic prediction Forecasting
[462] 2023 Health Epidemic prediction Forecasting
[498] 2023 Health Disease diagnosis Classification
[434] 2023 Health Clinical risk classification Classification
[270] 2024 Health Infectious disease Forecasting
[527] 2024 Health EEG emotion recognition Classification
[119] 2024 Health EEG emotion recognition Classification
[445] 2024 Health Epidemic Forecasting
[528] 2024 Health Epidemic Forecasting
[285] 2024 Health Epidemic Forecasting
[187] 2024 Health Epidemic Forecasting
[134] 2020 Mobility E-Scooter distribution Forecasting
[135] 2020 Mobility Bike sharing Forecasting
[393] 2020 Mobility Urban traffic Forecasting
[8] 2020 Mobility Urban traffic Forecasting
[33] 2020 Mobility Urban traffic Forecasting
[496] 2020 Mobility Urban traffic Forecasting
[275] 2020 Mobility Urban traffic Forecasting
[308] 2020 Mobility Urban traffic Forecasting
[301] 2020 Mobility Metro flow Forecasting
[431] 2020 Mobility Urban traffic Forecasting
[39] 2020 Mobility Urban traffic Forecasting